OpenStudy (anonymous):

Which ordered pair is a solution of the equation 2x − y = 9?

OpenStudy (anonymous):

a. -4,1 b. -2,5 c. 5,1 d. 6,-3

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

2(-4)-1=9 -8-1=9 -10=9 (NO) 2(-2)-5=9 -4-5=9 -9=9 (NO) 2(5)-1=9 10-1=9 9=9 (YES) 2(6)-(-3)=9 12+3=9 15=9 (NO)

OpenStudy (triciaal):



OpenStudy (anonymous):

Naah we do it my way and my way is easier

OpenStudy (anonymous):

Right @Kaelyn78?

OpenStudy (anonymous):

Yes, right.